OpenStudy (anonymous):

One plant from a garden is randomly selected. There is a 30% chance that the plant was purchased this year, and there is a 6% chance that the plant was purchased this year and produces berries. What is the probability that the chosen plant does not produce berries, given that the plant was purchased this year? Write the probability as a percent. Round to the nearest tenth if needed.

OpenStudy (anonymous):

@uri

OpenStudy (mathmate):

 One way to look at it is by the Venn Diagram. 30 out of 100 was purchased this year. 6 of those (from this year, namely 30) produced berries. How many of those from this year did not produce berries? What is the corresponding probability (out of the 30 from this year) that plants did not produce berries?

OpenStudy (anonymous):

THanks for doing the step by step

OpenStudy (mathmate):

Alternatively, use definition of conditional probability: P(no berries | from this year)=P(no berries and plants from this year)/P(plants from this year) =(24/100)/(30/100)

OpenStudy (anonymous):

8000?

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

or 0.8

OpenStudy (mathmate):

Probabilities are always between 0 and 1. If it goes outside, check and double check. Hope you put in the parentheses when you evaluated the expression above.

OpenStudy (mathmate):

Yes, 0.8 is more like it. But do you understand how it works?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

is this the right answer

OpenStudy (mathmate):

 If we're given that the plants are from this year, the sample space is 30, so 24/30=0.8.